Looking for some help. this question is a smoker.
Among 100 people randomly chosen in a survey, 10 stated that they believed that the elected representatives are honest most of the time. The approximate 95% confidence interval for the percentage of the population who believe that such people are honest most of the time is
thank you for your time and help.

Answer by stanbon(75887)

Among 100 people randomly chosen in a survey, 10 stated that they believed that the elected representatives are honest most of the time. The approximate 95% confidence interval for the percentage of the population who believe that such people are honest most of the time is
----------
p-hat = 10/100=0.1
E=z*sqrt(pq/n)
E=1.96sqrt(0.1*0.9/100)
E=1.96*0.0009
E=0.0018
CI: (p-hat-E,p-hat+E)
CI: (0.1-0.0018, 0.1+0.0018)
CI: (0.0982,0.1018)
============
Cheers,
Stan H.